OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (johnweldon1993):

Lets do the substitution method... Solve the second equation for (lets do 'x') Add 3y to both sides 2x - 3y = 10 +3y +3y ---------------- 2x = 3y + 10 <---now divide everything by 2 to solve for 'x' ------------ 2 2 2 x = (3/2)y + 5 Now that we know what 'x' equals...we substitute that into the first equation we have for 'x' 3x + 4y = -2 After substituting that in...we have \[\large 3(\frac{3}{2}y + 5) + 4y = -2 \] Simpify a bit... \[\large \frac{9}{2}y + 15 + 4y = -2\] Combine like terms... \[\large \frac{17y}{2} + 15 = -2\] Subtract 15 from both sides... \[\large \frac{17y}{2} = -17\] Multiply both sides by 2 \[\large 17y = -34\] Divide both sides by 17 \[\large y = -2\] Now that we know y = -2 we can use that to solve for 'x' go back to our first original equation... \[\large 3x + 4y = -2\] We know y = -2 so sub that in... \[\large 3x + 4(-2) = -2\] \[\large3x - 8 = -2\] Add 8 to both sides... \[\large 2x = 6\] Divide both sides by 2 \[\large x = 3\] Now we solved and the answer is x = 3 and y = -2

OpenStudy (johnweldon1993):

No no no hang on a second

OpenStudy (johnweldon1993):

made a slight error in the LAST bit of that... After add 8 to both sides..it should be \[\large 3x = 6\] Then divide both sides by 3 not 2 \[\large x =2 \] So x = 2 and y = -2

OpenStudy (anonymous):

ahhh thank you so much for working it out step by step it helped me understand a lot @johnweldon1993

OpenStudy (johnweldon1993):

No problem @jadeedwardson :) if you need anymore help feel free to message me :)